OpenStudy (anonymous):

if arc AB measures 155 degrees (14 cm), what is the length of the radius of the circle, to the nearest hundredth? drawing.........

OpenStudy (anonymous):



OpenStudy (anonymous):

my teacher says the answer is 5.71 (the radius)

OpenStudy (anonymous):

...i just dont know how to solve it

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Arc_(geometry) \[\frac{arc degree}{360}=\frac{arclength}{circumference}\] http://en.wikipedia.org/wiki/Arc_(geometry) \[\frac{arc degree}{360}=\frac{arclength}{circumference}\] \[\frac{155}{360}=\frac{14}{2\pi*r}\] Using cross multiplication: \[155*2\pi*r=14*360\] Solve for r. Note: I got 5.17, so someone wrote it down wrong (either you or the teacher!)

OpenStudy (anonymous):

lol i was just going to ask if i could cross multiply

OpenStudy (anonymous):

omg thanks.